# 3]Sorbonne Université, CNRS, Université de Paris, Inria, Laboratoire Jacques-Louis Lions UMR 7598, 75005 Paris, France

### v1

@@ 3.1 Diffusion-based movement models
Here the vector ${\bm{S}}(t,{\bm{x}})=\left(S_{1}(t,{\bm{x}}),\ldots,S_{N}(t,{\bm{x}})\right)$ represents the concentrations at position ${\bm{x}}$ at time $t$ of $N$ different abiotic factors, whose dynamics are governed by appropriate evolution equations coupled with the PS-PIDE (13). In this more general context, definitions (11)-(12) can be modified, respectively, as
@@
Starting from [164], where a PS-PIDE model of form (13) with $D({\bm{y}})\equiv 0$ and $\bar{D}=0$ was proposed to study the emergence of cell phenotypic heterogeneity in avascular tumours, models that can be regarded as variations on the theme of (13) have found application in a number of theoretical studies on the eco-evolutionary dynamics of cancer cells in avascular and vascularised tumours – e.g. see [48, 53, 52, 93, 176, 237, 238]. Variants of (13) with $D({\bm{y}})\equiv D$ have also received attention from the mathematical community as a mean to investigate analytically evolutionary dynamics in phenotype-structured populations exposed to shifting environments (Alfaro et al. 5, 2, Berestycki and Chapuisat 25) and spatially periodic environments [44].
@@ 3.2 Pressure-based movement models
Figure 6: Illustration of phenotypic structuring across travelling waves from a PS-PIDE model of pressure-based movement. Illustrative outputs of (b) a PS-PIDE model of pressure-based movement of form (17) and (a) a corresponding ABM of the type presented in Section 4.1.1. In particular, phenotypic structuring across travelling waves is shown for the rescaled PS-PIDE (57) with $\varepsilon=0.01$ and the corresponding rescaled ABM. Here, we consider one dimension in each of physical and phenotype space, such that ${\bm{x}}\equiv x\in[0,25]$ and ${\bm{y}}\equiv y\in[0,1]$. Zero-flux conditions are imposed at the boundaries. These results are for a go-or-grow type scenario, where the phenotypic state $y\in[0,1]$ is such that $y$ close to 00 corresponds to high proliferation rate and low mobility coefficient, while $y$ close to 1 corresponds to low proliferation rate and high mobility coefficient. We see an invasion process in which a travelling wave of cells progresses, and there is a good agreement between the results of numerical simulations of the ABM, numerical solutions of the PS-PIDE model, and analytical predictions (right panels); cells with a higher mobility coefficient are concentrated to the front of the wave, and cells with a higher proliferation rate dominate the rear (left panels). The insets of the right panels display the plots of the points $x_{1}(t)$, $x_{2}(t)$, and $x_{3}(t)$ such that $\rho(t,x_{1}(t))=0.2\rho_{M}$, $\rho(t,x_{2}(t))=0.5\rho_{M}$, and $\rho(t,x_{3}(t))=0.8\rho_{M}$, where $\rho_{M}\in\mathbb{R}^{+}_{*}$ is the maximum value of the cell density. Notably, we observe the evolution of $x_{1}(t)$, $x_{2}(t)$, and $x_{3}(t)$ towards straight lines of approximately the same slope, and the slope agrees with analytical predictions on the wave speed. This indicates that there is evolution towards a travelling wave with a wave speed consistent with analytical predictions. For full details of the models and parametrisation, we refer to [189]: this figure corresponds to Figure 2 of [189].
@@
Diffusion-based models of cell movement implicitly rely on the assumption that cell motion is undirected. However, cells, especially eukaryotic cells, can have a tendency to move towards regions where they feel less compressed [40, 24]. This aspect of cell movement, which is central to a variety of physiological and pathological processes – encompassing tissue development, wound healing, and avascular tumour growth – can be captured by pressure-based models [212]. Building on the ideas presented in the seminal paper by [120] and subsequent papers [14, 27, 31, 40, 24, 51, 160, 209, 217], the basic tenet of these models is that cell movement is described via an advective term with an advective velocity inversely proportional to the gradient of the cellular pressure, $P(t,{\bm{x}})$. This leads to a PDE of the following form, which can be obtained from (1) by taking $D_{\bm{x}}\equiv 0$, ${\bm{A}_{\bm{x}}}:=-\mu\nabla_{\bm{x}}P$, and $F:=\rho R(P)$:
@@
The growth rate of the cell density, $R$, depends on $P$ to incorporate pressure-dependent inhibition of growth (i.e. cessation of cell division when cellular pressure becomes too high) [73, 211]. Note that, if the growth of the cell population is not only pressure-regulated but also nutrient-limited, then the growth rate would also depend on the nutrient concentration, and an additional PDE for the nutrient concentration would be introduced. Models of form (16) – and related variants for multiple cell populations – have drawn interest from mathematicians and physicists for their ability to recapitulate key aspects of tumour and tissue growth and for exhibiting travelling waves which display interesting features – e.g. see [34, 64, 166, 234].
Note also that when one takes $\Pi(\rho):=K\rho$, where $K\in\mathbb{R}^{+}_{*}$ is a scale factor (which for simplicity can be set as $1$), the transport term in the PDE (16) models more specifically the tendency of cells to disperse to avoid crowding (i.e. to move down the gradient of the cell density $\rho$[54], while the reaction term takes into account density-dependent inhibition of growth [161]. Moreover, as discussed in [208], when $\Pi(\rho):=K_{\gamma}\big{(}\rho\big{)}^{\gamma}$, where the parameter $\gamma\in\mathbb{R}^{+}_{*}$ with $\gamma>1$ provides a measure of the stiffness of the pressure law and $K_{\gamma}\in\mathbb{R}^{+}_{*}$ is a scale factor, the PDE (16) takes the form of a porous medium-type equation. In this case, considering the asymptotic regime $\gamma\to\infty$ [^14] Since the parameter $\gamma$ provides a measure of the stiffness of the pressure law $\Pi(\rho)$, the limiting regime $\gamma\to\infty$ is usually referred to as ‘incompressible limit’, because it corresponds to mathematically approximating cells as an incompressible fluid., from the PDE (16) one can derive a free-boundary problem of Hele-Shaw type, as demonstrated in [208], thus forming a bridge between PDE models of form (16) and mathematical models formulated as free-boundary problems, which have been widely employed to describe related biological processes [106]. This is also an aspect that has received increasing attention from the mathematical community – see e.g. [43, 79, 147, 187].
Generalisations of PDE models of form (16) into PS-PIDE models of form
@@
have been proposed for the situation in which both the rate of proliferation and the mobility coefficient of the cells depend on the phenotypic state ${\bm{y}}$. Specifically, a PS-PIDE model of this form with $\Pi(\rho):=\rho$ has been proposed in [173], while different possible definitions of $\Pi(\rho)$ have been considered in [189]. PS-PIDE models of type (17) can be obtained from (2) by taking $D_{\bm{x}}\equiv 0$ and ${\bm{A}_{\bm{x}}}:=-\mu({\bm{y}})\nabla_{\bm{x}}P$ along with $F:=nR({\bm{y}},P)$, ${\bm{A}}_{\bm{y}}\equiv 0$, and $D_{\bm{y}}\equiv\bar{D}$. Here the function $\mu({\bm{y}})$ is the mobility coefficient of cells in the phenotypic state ${\bm{y}}$. Furthermore, the function $R({\bm{y}},P)$ models the net proliferation rate of cells in the phenotypic state ${\bm{y}}$ under the local environmental conditions corresponding to the cellular pressure $P$, and can thus be regarded as a fitness function. This can be defined along the lines of (11) as
@@
where the parameter $\kappa\in\mathbb{R}^{+}_{*}$ is now linked to the critical value of the cellular pressure above which cessation of cell division occurs.
In [173, 189], considering ‘go-or-grow’ type scenarios in which fast proliferating cells are less migratory (i.e. they have a lower mobility coefficient) and vice versa, it was shown that PS-PIDEs of form (17) can produce phenotype-structured travelling waves whereby fast proliferating cells make up the bulk of the population in the rear of the wave, while highly migratory cells drive invasion at the edge of the wave front (see Figure 6). Focusing on the case where all cells in the population have the same mobility coefficient regardless of their phenotypic state (i.e. when $\mu({\bm{y}})\equiv\mu$), connections between PS-PIDE models of this type and free boundary problems of Hele-Shaw type have been explored in [70], while related models have been employed to investigate the emergence of resistance to chemotherapy alongside tumour growth in [59]. Moreover, extensions of (17) have been considered in [92] to model the growth of avascular tumours, taking into account tumour necrosis and tumour-microenvironment interactions.
@@ 3.3 Taxis-based movement models
Figure 7: Illustration of phenotypic structuring across travelling waves from a PS-PIDE model of chemotaxis-based movement. Illustrative outputs of a PS-PIDE model of chemotaxis-based movement of form (20). In particular, phenotypic structuring across travelling waves is shown for the rescaled PS-PIDE  (56) with $\varepsilon=0.01$ coupled with an appropriate evolution equation for the attractant concentration, $S$. Here, we consider one dimension in each of physical and phenotype space, such that ${\bm{x}}\equiv x\in[0,20]$ and ${\bm{y}}\equiv y\in[0,1]$. Zero-flux conditions are imposed at the boundaries. (a) Chemotaxis-proliferation trade-off, where the phenotypic state $y\in[0,1]$ is such that $y$ close to 00 corresponds to high proliferation rate and low chemotactic sensitivity, while $y$ close to 1 corresponds to low proliferation rate and high chemotactic sensitivity. All cells equally degrade the attractant. We see an invasion process in which as the wave progresses, nutrient is degraded (left panel); the most efficient chemotaxers are concentrated to the front of the wave, and strong proliferators dominate the rear (middle panel); the numerically estimated wavespeed agrees with analytical predictions (right panel). (b) Chemotaxis-proliferation trade-off for a nutrient-like attractant, i.e. an attractant that impacts on the proliferation rate. The phenotypic state $y\in[0,1]$ is again such that $y$ close to 00 corresponds to high proliferation rate but low chemotactic sensitivity, while $y$ close to 1 corresponds to low proliferation rate and high chemotactic sensitivity. We see a pulse waveform for the population (left panel); as in (a), strongly (weakly) chemotactic cells are concentrated to the front (rear) of the wave (middle panel); numerically estimated wavespeed agrees with analytical predictions (right panel). For full details of the model and parametrisation, we refer to [173]: panel (a) corresponds to Figure 3A and panel (b) corresponds to Figure 5B of [173], respectively.
@@
In processes that include morphogenesis, wound healing, and inflammation, cell populations need to be correctly positioned. Taxis-based [^15] The word taxis stems from the ancient Greek $\tau\alpha\xi\iota\sigma$, meaning ‘to steer’, and was adopted by Wiliam Pfeffer to describe a hypothetical steering of microorganisms, following observations of their reorganisation according to chemicals and light [205]. models form a third major class of movement models, describing situations in which a cell (or organism) is directed to move in a particular direction. Chemotaxis, the directed movement of cells up or down gradients of (usually soluble) molecules, has been the most intensely studied form of taxis behaviour, but a wide variety of others have also been investigated, including movement in response to adhesion gradients (haptotaxis, Carter 46), stiffness (durotaxis, Lo et al. 177), and electric fields (galvanotaxis, Erickson and Nuccitelli 87).
@@
The most well-known PDE model for describing chemotaxis movements is the Keller-Segel system [149, 151, 150, 199]. This model has received significant attention as a model for chemotaxis [130, 229, 131, 23, 198, 17]: from the modelling community for its ability to recapitulate macroscopic dynamics of chemotactic populations, including travelling waves and self-organisation, and from the mathematical community for its intricate analytical properties. Moreover, it has provided an important base for describing other forms of taxis behaviour, with its adaption to describe haptotaxis a particularly prominent example [246, 218]. In its standard form, (chemo)taxis is described via an advective term with an advective velocity proportional to the gradient of a (chemical) signal, $S(t,{\bm{x}})$. Along with standard choices of an additional linear diffusion component (i.e. similarly to what done in (5)) and a growth term, one sets $D_{\bm{x}}\equiv D$, ${\bm{A}_{\bm{x}}}:=\chi\nabla_{\bm{x}}S$, and $F:=\rho R(\rho,S)$ in (1) to obtain
@@
An extension of (19) to a PS-PIDE model for chemotaxis has been proposed in [172], focusing on the case where phenotype regulates both proliferation and chemotaxis. One can arrive at this form of model through extending to a phenotype density $n(t,{\bm{x}},{\bm{y}})$ and – with the same earlier choice of linear diffusion in phenotype (i.e. taking ${\bm{A}}_{\bm{y}}\equiv 0$ and $D_{\bm{y}}\equiv\bar{D}$ in (2)) – obtaining the following PS-PIDE formulation of (19):
@@
In the above model, each of the chemotactic sensitivity and growth rate now depend on the phenotypic state – i.e. the function $\chi({\bm{y}})$ models the chemotactic sensitivity of cells in the phenotypic state ${\bm{y}}$, while the function $R({\bm{y}},\rho,S)$ is the net proliferation rate (i.e. the fitness) of cells in the phenotypic state ${\bm{y}}$ under the local environmental conditions corresponding to the cell density $\rho$ and the signal $S$, which can for example be defined via (14) with ${\bm{S}}\equiv S$. Experimental studies with microfluidic T-mazes have revealed that bacterial populations can be structured according to their chemotactic sensitivity (Salek et al. 219, see Figure 2 (b)), while other studies have suggested that resource allocation can lead to a trade-off between investment in motility and growth [193]. Consequently, the case of a ‘proliferation-chemotaxis’ trade-off – that is, faster proliferators are less chemotactic (i.e. they have a lower chemotactic sensitivity) and vice versa – was considered in [172], where it was demonstrated that PS-PIDEs of form (20) coupled with an appropriate evolution equation for $S$ can form travelling waves that exhibit phenotypic structuring (see Figure 7).
@@
Of course, one can consider a number of further variations. For instance, letting $S$ model the density of the extracellular matrix, PS-PIDEs of form (20) have been employed to describe cancer cell dynamics in phenotype-structured haptotaxis models of tumour invasion [92, 168]. Related PS-PIDE models of cell migration, where $S$ is replaced by a vector ${\bm{S}}$ whose components model the concentrations of different diffusible molecules, have been considered in [133]. Variants of PS-PIDEs of form (20), wherein a phenotypic structure is included in the growth term but not in the term modelling cellular taxis, have been used in models for cell migration [81] and tumour cell invasion [90, 125, 126].
@@
PS-PIDE models closely related to (20) have also been elegantly deployed alongside experiments, to investigate how phenotype structuring drives collective behaviour within travelling E. coli bands [98, 182, 207]. In these models, the structuring describes a bacterium’s tumbling bias and manifests in phenotype dependency for both cell diffusion, $D_{\bm{x}}:=D({\bm{y}})$, and chemotactic sensitivity, $\chi({\bm{y}})$. These models assume that bacteria do not alter phenotype within their lifetime (i.e. ${\bm{A}}_{\bm{y}}\equiv 0$ and $D_{\bm{y}}\equiv 0$ in (2)), but can as a result of division, an aspect that is included in the model via a kernel that stipulates how division of a parent cell with one phenotype leads to progeny cells with other phenotypes. This leads to a non-local form of the term $F$ in (2) that depends on the phenotype structuring. Direct comparisons of this model with data have revealed the limitations of ‘traditional’ non-structured models and gained insights into how an evolving structuring may allow populations to efficiently colonise heterogeneous environments.

### v2

@@ 3.1 Diffusion-based movement models
Here the vector ${\bm{S}}(t,{\bm{x}})=\left(S_{1}(t,{\bm{x}}),\ldots,S_{N}(t,{\bm{x}})\right)$ represents the concentrations at position ${\bm{x}}$ at time $t$ of $N$ different abiotic factors, the dynamics of which are governed by appropriate evolution equations coupled with the PS-PDE (13). In this more general context, definition (11) and definition (12) can be modified, respectively, as
@@
Starting from [176], where a PS-PDE model of form (13) with $D({\bm{y}})\equiv 0$ and $\bar{D}=0$ was proposed to study the emergence of cell phenotypic heterogeneity in avascular tumours, models that can be regarded as variations on the theme of (13) have found application in a number of theoretical studies on the eco-evolutionary dynamics of cancer cells in avascular and vascularised tumours – e.g. see [51, 58, 57, 94, 170, 245, 244]. Variants of (13) with $D({\bm{y}})\equiv D$ have also received attention from the mathematical community as a mean to investigate analytically evolutionary dynamics in phenotype-structured populations exposed to shifting environments (Alfaro et al 6, 7, Berestycki and Chapuisat 27) and spatially periodic environments [38].
@@ 3.2 Pressure-based movement models
Figure 6: Illustration of phenotypic structuring across travelling waves from a PS-PDE model of pressure-based movement. Illustrative outputs of (b) a PS-PDE model of pressure-based movement of form (17) and (a) a corresponding ABM of the type presented in Section 4.1.1. In particular, phenotypic structuring across travelling waves is shown for the rescaled PS-PDE (57) with $\varepsilon=0.01$ and the corresponding rescaled ABM. Here, we consider one dimension in each of physical and phenotype space, such that ${\bm{x}}\equiv x\in[0,25]$ and ${\bm{y}}\equiv y\in[0,1]$. Zero-flux conditions are imposed at the boundaries. These results are for a go-or-grow type scenario, where the phenotypic state $y\in[0,1]$ is such that $y$ close to 00 corresponds to high proliferation rate and low mobility coefficient, while $y$ close to 1 corresponds to low proliferation rate and high mobility coefficient. We see an invasion process in which a travelling wave of cells progresses, and there is a good agreement between the results of numerical simulations of the ABM, numerical solutions of the PS-PDE model, and analytical predictions (right panels). Cells with a higher mobility coefficient are concentrated to the front of the wave, and cells with a higher proliferation rate dominate the rear (left panels). The insets of the right panels display the plots of the points $x_{1}(t)$, $x_{2}(t)$, and $x_{3}(t)$ such that $\rho(t,x_{1}(t))=0.2\rho_{M}$, $\rho(t,x_{2}(t))=0.5\rho_{M}$, and $\rho(t,x_{3}(t))=0.8\rho_{M}$, where $\rho_{M}\in\mathbb{R}^{+}_{*}$ is the maximum value of the cell density. Notably, we observe the evolution of $x_{1}(t)$, $x_{2}(t)$, and $x_{3}(t)$ towards straight lines of approximately the same slope, and the slope agrees with analytical predictions on the wave speed. This indicates that there is evolution towards a travelling wave with a wave speed consistent with analytical predictions. For full details of the models and parametrisation, we refer to [182]: this figure corresponds to Figure 2 of [182].
@@
Diffusion-based models of cell movement implicitly rely on the assumption that cell motion is undirected. However, cells, especially eukaryotic cells, can have a tendency to move towards regions in which they are less compressed [45, 43]. This aspect of cell movement, which is central to a variety of physiological and pathological processes – encompassing tissue development, wound healing, and avascular tumour growth – can be captured by pressure-based models [216]. Building on the ideas presented in the seminal paper by [116] and subsequent papers [13, 39, 44, 45, 43, 64, 180, 213, 224], the basic tenet of these models is that cell movement is described via an advective term with an advective velocity inversely proportional to the gradient of the cellular pressure, $P(t,{\bm{x}})$. This leads to a PDE of the following form, which can be obtained from (1) by taking $D_{\bm{x}}\equiv 0$, ${\bm{A}_{\bm{x}}}:=-\mu\nabla_{\bm{x}}P$, and $F:=\rho R(P)$:
@@
The growth rate of the cell density, $R$, depends on $P$ to incorporate pressure-dependent inhibition of growth (i.e. cessation of cell division when cellular pressure becomes too high) [81, 214]. Note that, if the growth of the cell population was not only pressure-regulated but also nutrient-limited, then the growth rate would also depend on the nutrient concentration, and an additional PDE for the nutrient concentration would be introduced. Models of form (16) – and related variants for multiple cell populations – have drawn interest from mathematicians and physicists for their ability to recapitulate key aspects of tumour and tissue growth, and also for exhibiting travelling waves which display interesting features – e.g. see [31, 47, 54, 169, 231].
Note also that when one takes $\Pi(\rho):=K\rho$, where $K\in\mathbb{R}^{+}_{*}$ is a scale factor (which for simplicity can be set as $1$), the transport term in the PDE (16) models more specifically the tendency of cells to disperse to avoid crowding (i.e. to move down the gradient of the cell density $\rho$[53], while the reaction term takes into account density-dependent inhibition of growth [161]. Moreover, as discussed by [209], when $\Pi(\rho):=K_{\gamma}\rho^{\gamma}$, where the parameter $\gamma\in\mathbb{R}^{+}_{*}$ with $\gamma>1$ provides a measure of the stiffness of the pressure law and $K_{\gamma}\in\mathbb{R}^{+}_{*}$ is a scale factor, the PDE (16) takes the form of a porous medium-type equation. In this case, considering the asymptotic regime $\gamma\to\infty$ [^15] Since the parameter $\gamma$ provides a measure of the stiffness of the pressure law $\Pi(\rho)$, the limiting regime $\gamma\to\infty$ is usually referred to as ‘incompressible limit’, because it corresponds to mathematically approximating cells as an incompressible fluid., from the PDE (16) one can derive a free-boundary problem of Hele-Shaw type, as demonstrated by [209]. This forms a bridge between PDE models of form (16) and mathematical models formulated as free-boundary problems, which have also been widely employed [110]. This is also an aspect that has received increasing attention from the mathematical community – see e.g. [41, 73, 150, 187].
Generalisations of PDE models of the form (16) into PS-PDE models of form
@@
have been proposed for the situation in which both the rate of proliferation and the mobility coefficient of the cells depend on the phenotypic state ${\bm{y}}$. Specifically, a PS-PDE model of this form with $\Pi(\rho):=\rho$ has been proposed by [171], while different possible definitions of $\Pi(\rho)$ have been considered by [182]. PS-PDE models of type (17) can be obtained from (2) by taking $D_{\bm{x}}\equiv 0$ and ${\bm{A}_{\bm{x}}}:=-\mu({\bm{y}})\nabla_{\bm{x}}P$ along with $F:=nR({\bm{y}},P)$, ${\bm{A}}_{\bm{y}}\equiv 0$, and $D_{\bm{y}}\equiv\bar{D}$. Here the function $\mu({\bm{y}})$ is the mobility coefficient of cells in the phenotypic state ${\bm{y}}$. Furthermore, the function $R({\bm{y}},P)$ models the net proliferation rate of cells in the phenotypic state ${\bm{y}}$ under the local environmental conditions corresponding to the cellular pressure $P$, and can thus be regarded again as a fitness function. This can be defined along the lines of (11) as
@@
where the parameter $\kappa\in\mathbb{R}^{+}_{*}$ is inversely related to the critical value of the cellular pressure above which cessation of cell division occurs.
Considering ‘go-or-grow’ type scenarios in which fast proliferating cells are less migratory (i.e. they have a lower mobility coefficient) and vice versa, [171, 182] have shown that PS-PDEs of form (17) can produce phenotype-structured travelling waves whereby fast proliferating cells make up the bulk of the population in the rear of the wave, while highly migratory cells drive invasion at the edge of the wave front (see Figure 6). Focusing on the case where all cells in the population have the same mobility coefficient regardless of their phenotypic state (i.e. when $\mu({\bm{y}})\equiv\mu$), connections between PS-PDE models of this type and free boundary problems of Hele-Shaw type have been explored by [71], while related models have been employed to investigate the emergence of resistance to chemotherapy alongside tumour growth by [63]. Moreover, extensions of (17) have been considered by [95] to model the growth of avascular tumours, taking into account tumour necrosis and tumour-microenvironment interactions.
@@ 3.3 Taxis-based movement models
Figure 7: Illustration of phenotypic structuring across travelling waves from a PS-PDE model of chemotaxis-based movement. Illustrative outputs of a PS-PDE model of chemotaxis-based movement of form (20). In particular, phenotypic structuring across travelling waves is shown for the rescaled PS-PDE  (56) with $\varepsilon=0.01$ coupled with an appropriate evolution equation for the attractant concentration, $S$. Here, we consider one dimension in each of physical and phenotype space, such that ${\bm{x}}\equiv x\in[0,20]$ and ${\bm{y}}\equiv y\in[0,1]$. Zero-flux conditions are imposed at the boundaries. (a) Chemotaxis-proliferation trade-off, where the phenotypic state $y\in[0,1]$ is such that $y$ close to 00 corresponds to high proliferation rate and low chemotactic sensitivity, while $y$ close to 1 corresponds to low proliferation rate and high chemotactic sensitivity. All cells equally degrade the attractant. We see an invasion process in which as the wave progresses, nutrient is degraded (left panel); the most efficient chemotaxers are concentrated to the front of the wave, and strong proliferators dominate the rear (middle panel); the numerically estimated wavespeed agrees with analytical predictions (right panel). (b) Chemotaxis-proliferation trade-off for a nutrient-like attractant, i.e. an attractant that has an impact on the cell proliferation rate. The phenotypic state $y\in[0,1]$ is again such that $y$ close to 00 corresponds to high proliferation rate but low chemotactic sensitivity, while $y$ close to 1 corresponds to low proliferation rate and high chemotactic sensitivity. We see a pulse waveform for the population (left panel); as in (a), strongly (weakly) chemotactic cells are concentrated to the front (rear) of the wave (middle panel); numerically estimated wavespeed agrees with analytical predictions (right panel). For full details of the model and parametrisation, we refer to [171]: panel (a) corresponds to Figure 3A and panel (b) corresponds to Figure 5B of [171], respectively.
@@
In processes that include morphogenesis, wound healing, and inflammation, cell populations need to be correctly positioned. Taxis-based [^16] The word taxis stems from the ancient Greek $\tau\alpha\xi\iota\sigma$, meaning ‘to steer’, and was adopted by Wiliam Pfeffer to describe a hypothetical steering of microorganisms, following observations of their reorganisation according to chemicals and light [210]. models form a third major class of movement models, describing situations in which a cell (or organism) is directed to move in a particular direction. Chemotaxis, the directed movement of cells up or down gradients of (usually soluble) molecules, has been the most intensely studied form of taxis behaviour, but a wide variety of others have also been investigated, including movement in response to adhesion gradients (haptotaxis, Carter 48), stiffness (durotaxis, Lo et al 164), and electric fields (galvanotaxis, Erickson and Nuccitelli 90).
@@
The most well-known PDE model for describing chemotactic movements is the Keller-Segel system [144, 146, 145, 204]. This model has received significant attention as a model for chemotaxis [133, 233, 128, 25, 202, 17]: from the modelling community for its ability to recapitulate macroscopic dynamics of chemotactic populations, including travelling waves and self-organisation, and from the mathematical community for its intricate analytical properties. Moreover, it has provided an important base for describing other forms of taxis behaviour, with its adaption to describe haptotaxis a particularly prominent example [248, 221]. In its standard form, (chemo)taxis is described via an advective term with an advective velocity proportional to the gradient of a (chemical) signal, $S(t,{\bm{x}})$. Along with standard choices of an additional linear diffusion component (i.e. similarly to what done in (5)) and a growth term, one sets $D_{\bm{x}}\equiv D$, ${\bm{A}_{\bm{x}}}:=\chi\nabla_{\bm{x}}S$, and $F:=\rho R(\rho,S)$ in (1) to obtain
@@
An extension of (19) to a PS-PDE model for chemotaxis has been proposed by [165], focusing on the case where phenotype regulates both proliferation and chemotaxis. One can arrive at this form of model through extending to a phenotype density $n(t,{\bm{x}},{\bm{y}})$ and – with the same earlier choice of linear diffusion in phenotype (i.e. taking ${\bm{A}}_{\bm{y}}\equiv 0$ and $D_{\bm{y}}\equiv\bar{D}$ in (2)) – obtaining the following PS-PDE formulation of (19):
@@
In the above model, each of the chemotactic sensitivity and growth rate now depend on the phenotypic state – i.e. the function $\chi({\bm{y}})$ models the chemotactic sensitivity of cells in the phenotypic state ${\bm{y}}$, while the function $R({\bm{y}},\rho,S)$ is the net proliferation rate (i.e. the fitness) of cells in the phenotypic state ${\bm{y}}$ under the local environmental conditions corresponding to the cell density $\rho$ and the signal $S$, which can for example be defined via (14) with ${\bm{S}}\equiv S$. Experimental studies with microfluidic T-mazes have revealed that bacterial populations can be structured according to their chemotactic sensitivity (Salek et al 218, see Figure 2 (b)), while other studies have suggested that resource allocation can lead to a trade-off between investment in motility and growth [196]. Consequently, the case of a ‘proliferation-chemotaxis’ trade-off – that is, faster proliferators are less chemotactic (i.e. they have a lower chemotactic sensitivity) and vice versa – was considered by [165], who demonstrated that PS-PDEs of form (20) coupled with an appropriate evolution equation for $S$ can form travelling waves that exhibit phenotypic structuring (see Figure 7).
@@
Of course, one can consider a number of further variations. For instance, letting $S$ model the density of the extracellular matrix, PS-PDEs of form (20) have been employed to describe cancer cell dynamics in phenotype-structured haptotaxis models of tumour invasion [95, 173]. Related PS-PDE models of cell migration, where $S$ is replaced by a vector ${\bm{S}}$, the components of which model the concentrations of different diffusible molecules, have been considered by [130]. Variants of PS-PDEs of form (20), wherein a phenotypic structure is included in the growth term but not in the term modelling cellular taxis, have been used in models for cell migration [79] and tumour cell invasion [88, 129, 131].
@@
PS-PDE models closely related to (20) have also been elegantly deployed alongside experiments, to investigate how phenotype structuring drives collective behaviour within travelling E. coli bands [111, 184, 212]. In these models, the structuring describes a bacterium’s tumbling bias and manifests in phenotype dependency for both cell diffusion, $D_{\bm{x}}:=D({\bm{y}})$, and chemotactic sensitivity, $\chi({\bm{y}})$. These models assume that bacteria do not alter phenotype within their lifetime (i.e. ${\bm{A}}_{\bm{y}}\equiv 0$ and $D_{\bm{y}}\equiv 0$ in (2)), but can as a result of division, an aspect that is included in the model via a kernel that stipulates how division of a parent cell with one phenotype leads to progeny cells with other phenotypes. This leads to a non-local form of the term $F$ in (2) that depends on the phenotype structuring. Direct comparisons of this model with data have revealed the limitations of ‘traditional’ non-structured models and gained insights into how an evolving structuring may allow populations to efficiently colonise heterogeneous environments.